A stone is thrown upwards with an accelaration of 10 m/sec.sq.If its velocity is 10m/sec, then find the distance and taken to reach the ground.


Please suggest the right method to do it


thanks

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear NEEL bhatt

if you will through a stone in upward direction with an acceleration 10 m/sec2 in upword direction then it will keep movinf in upword direction ,and it will never come aback to ground.

 

but if the direction of acceleration is vertically downward .

then find the time when its velocity become zero

V= U -gt

0 = 10 - 10*t

 t =1 sec

ans same time it will take to come back to ground

so total time = 2*1 =2 sec

and for hight

V2 =U2 -2gh

0 = 100 -2*10 h

h =5 m

so total distance  H =2*5 =10 m

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin